rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#6182 — FS#10088 — sip.ovh.net

Attached to Project— VoIP
Incident
All phones
CLOSED
100%
Thereafter activating the new FWSM card in one of the VoIP routers, the MGCP and SIP phones
on the old infrastructure with the domain sip.ovh.net couldn't be recorded.

The re-recordings are in progress.
Date:  Wednesday, 29 January 2014, 11:23AM
Reason for closing:  Done
Additional comments about closing:  No more anomalies.
Comment by OVH - Saturday, 25 January 2014, 03:39AM

SIP phones are up to 80%, the MGCP are going up at their turn.


Comment by OVH - Saturday, 25 January 2014, 07:40AM

The SIP are completely UP, the MGCP are much difficult to recover (15%), we are searching the origin of this latency.


Comment by OVH - Saturday, 25 January 2014, 22:14PM

We are carrying on the researches on the absence of some MGCP recovery.


Comment by OVH - Saturday, 25 January 2014, 22:15PM

Thereafter the IP switching (cf: http://status.ovh.net/?do=details&id=6174) the MGCP are recovering more rapidly (30%).


Comment by OVH - Saturday, 25 January 2014, 22:16PM

We have switched the main IP, 60% of MGCP phones are UP.


Comment by OVH - Saturday, 25 January 2014, 22:18PM

http://status.ovh.net/?do=details&id=6174
We have switched a pure MGCP IP on the new
firewall and this has fixed the issue.

We have therefore switched the main IP SIP + MGCP.
We are at 60% UP.

This is finally going through ..


Comment by OVH - Saturday, 25 January 2014, 22:18PM

90% of the MGCP are UP.


Comment by OVH - Saturday, 25 January 2014, 22:19PM

We are having an ARP issue between the firewall blocking
50% of the traffic. We have done the clear of the IP
and do notice that the rest of the lines are recovering.

97%


Comment by OVH - Saturday, 25 January 2014, 22:22PM

100% of the SIP and MGCP lines are UP. Eventually, we will set a tool
to optimise the diagnosis of eventual packet loss on the infra.


Comment by OVH - Saturday, 25 January 2014, 22:28PM

100%.

Thus, all is UP. Some MGCP phones have a bad displaying
which is not reflecting the reality. We are updating
the display status.

We are on the new firewall which has a very huge capacity.
We shouldn't have issues due to the firewall.

We won't perform any intervention before Wednesday.
Starting from Wednesday, we will work on setting the
redundancy on the new firewall. We hope to be able to
set it in the next weekend.

We do apologise for the duration of the intervention and
issues related to the old firewall which is now cut and
could be buried with no regret.

The c5a and c5b infrastructure wasn't impacted by this
intervention. Only c4


Comment by OVH - Sunday, 26 January 2014, 23:32PM

We are again having loss of connection on sip.ovh.net


Comment by OVH - Sunday, 26 January 2014, 23:33PM

All the phones are again connected,

(Cf. http://status.ovh.net/?do=details&id=6184)


Comment by OVH - Monday, 27 January 2014, 07:46AM

There's still a problem on the firewall
which doesn't allow establishing any new connections.


Comment by OVH - Monday, 27 January 2014, 17:38PM

We will adapt the firewall configuration tonight
to correct this problem.


Comment by OVH - Monday, 27 January 2014, 22:53PM

The issue has reoccurred. We are applying the
modification immediately.


Comment by OVH - Monday, 27 January 2014, 22:54PM

Here we start.


Comment by OVH - Tuesday, 28 January 2014, 13:08PM

We have received feedback relating to a failure to retrieve calls
on hold and supervision on sip.ovh.net, an analysis is underway.


Comment by OVH - Tuesday, 28 January 2014, 16:11PM

According to our first analyses, it's related to the outbound proxy configuration
of the telephone on sip.ovh.fr, for domains already configured on sip.ovh.net,
SIP customers are no longer registered in this case.

We are continuing the tests with a particular case. Domain migration to sip.ovh.fr
or deactivation of the outbound proxy will resolve this issue.


Comment by OVH - Tuesday, 28 January 2014, 17:20PM

We have modified the configuration on our proxy SIP, 1000 SIP lines
were affected and have now been reset.

We are checking that everything is ok.


Comment by OVH - Wednesday, 29 January 2014, 11:23AM

We are no longer detecting any anomalies.